Using traditional methods, it takes 10.0 hours to receive a basic flying license. A new license training method using Computer Aided Instruction (CAI) has been proposed. A researcher used the technique with 16 students and observed that they had a mean of 10.2 hours with a standard deviation of 1.6. A level of significance of 0.05 will be used to determine if the technique performs differently than the traditional method. Assume the population distribution is approximately normal. Find the value of the test statistic. Round your answer to three decimal places.
Answer #1

To Test :-

H0 :-  \mu = 10

H1 :-  μメ10

Test Statistic :-
t = ( \bar{X} - \mu ) / (S /\sqrt{n})
(10.2-10)// (1.6N16) t
t = 0.500


Test Criteria :-
Reject null hypothesis if | t |\; > \;t_{\alpha /2, n-1}
to/2n-1 = t0.05/ 2.16-1-2.131
t! > taZn-1 = 0.5 < 2.131
Result :- Fail to reject null hypothesis


Decision based on P value
P - value = P ( t > 0.5 ) = 0.6243
Reject null hypothesis if P value < a-0.05 level of significance
P - value = 0.6243 > 0.05 ,hence we fail to reject null hypothesis
Conclusion :- Fail to reject null hypothesis

There is insufficient evidence to support the claim that the technique performs differently than the traditional method.